Given that log (lower a)(x)=3.58 and log (lower a) (y)= 4.79, find log (lower a)(y/x).
I can not figure out these logarithms. Please help. Test coming soon.

Answer by stanbon(48535)

Given that log (lower a)(x)=3.58 and log (lower a) (y)= 4.79, find log (lower a)(y/x).
Log Law: log(a/b)=loga - logb
Your problem
log(y/x)=logy - logx
=4.79 - 3.58
=1.21
Cheers,
Stan H.